Firm Z expects to earn $3 per share next year. the firm's ROE is 11% and its plowback ratio is 55%. if the firm's market capitalization rate is 8%, what is the present value growth opportunities?
Answer #1

Expected Earning, E1 = $3.00
Plowback Ratio, b = 55%

Payout Ratio = 1 - b
Payout Ratio = 1 - 0.55
Payout Ratio = 0.45

Expected Dividend, D1 = E1 * Payout Ratio
Expected Dividend, D1 = $3.00 * 0.45
Expected Dividend, D1 = $1.35

Growth Rate, g = ROE * b
Growth Rate, g = 11% * 0.55
Growth Rate, g = 6.05%

Market Capitalization Rate, k = 8.00%

Current Price, P0 = D1 / (k - g)
Current Price, P0 = $1.35 / (0.08 - 0.0605)
Current Price, P0 = $69.23

Present Value of Growth Opportunity, PVGO = P0 - (E1 / k)
Present Value of Growth Opportunity, PVGO = $69.23 - ($3.00 / 0.08)
Present Value of Growth Opportunity, PVGO = $69.23 - $37.50
Present Value of Growth Opportunity, PVGO = $31.73